Find values for m and b in the following system so that the solution to the system is (-3, 4).
5x+7y = b
mx+y = 22

Answer by malakumar_kos@yahoo.com(315)

Find values for m and b in the following system so that the solution to the system is (-3, 4).
5x+7y = b
mx+y = 22
substituting the values of x and y as -3 and 4 in the eq'ns we get
5(-3)+7(4) = b i.e -15+28 = b or b = 13
m(-3)+(4) = 22 i.e -3m = 22-4
-3m = 18
m = 18/-3 = -6
the values of m = -6 and b = 13